2017/3/28(Wed.)

 * Remote participants by eZuce: None

==== Report of work (*:responsible person) ====

 * Cry and Type A at ETMX: *Takahashi, Ishizaki, Sato, Okutomi, Tomaru, Takahashi, Ushiba, Miyamoto, Enomoto
  * Height was adjusted, and started yaw alignment. GAS filter had a sudden jump, and it is being fixed today.
  * First contact will be removed.
 * BS: *Enomoto
  * DAC had some glitches with no output (even dac kill and master SW are off).
  * It will be investigated more.

 * BRT: Akutsu
  * Alignment was done and BRT worked well, it means that the beam position can be controlled after BRT.
  * Beam shape looks elliptic a bit.


==== Work plan (*:responsible person) ====


==== Laser hazard ====
 * [AM]
  * IR: PSL, Pend 1F
  * Green: PR2
 * [PM]
  * IR: SR2-3 line, Xend 1F, Yend 1F
  * Green: PR2, Xend 1F 

 * [night]
  * IR: SR2-3 line, Xend 1F, Yend 1F
  * Green: PR2, Xend 1F 

==== Weekly ====


 * week of 3/26
  * 3/26 Pcal will start, and all the 1st floor of Y-end will be laser hazard area.
  * Connection of cable at ETMX and starting control.
  * Rough alignment of ETMX using IR.
   * 3/27 EX PD will be removed.: *Yokozawa

 * week of 4/2
  * ETMY damping.
  * Rough alignment of ETMX using IR, and control MICH if can.

 * week of 4/9


 * Issues

 * ETMX Type A: GAS filter had a sudden jump, and it is being fixed today.
 * BS: DAC had some glitches with no output (even dac kill and master SW are off).